OpenStudy (sonyalee77):

You began membership in a new health and fitness club which has access to a dietician and personal trainer. They help you develop a special eight-week diet and exercise program. The data in the following table represents your weight, w, as a function of time, t, over an eight-week period. Time (weeks) 0 1 2 3 4 5 6 7 8 Weight (lb) 150 144 143 141 140 137 137 132 129 If you lost 13 pounds in the first five weeks, what is the average rate of change of weight with respect to time over the first five weeks of the program? a. +0.38 pounds per week b. -0.38 pounds per week c. +2

OpenStudy (sonyalee77):

@Hero @Euler271

OpenStudy (anonymous):

weird question. logically, the only answer that can make sense is a negative one since it is a decreasing slope. the secant is the average rate of change (and is 1/tangent) the tangent is the instantaneous rate of change. (1/secant) we want the first one which is:\[\frac{ \Delta x }{ \Delta y } =\frac{ \Delta\ weeks }{ \Delta\ lbs }= \frac{ 5 - 0 }{ 137-150 } = \frac{ 5 }{ -13 } = -0.3846\]